# How Hungry is AI? Benchmarking Energy, Water, and Carbon Footprint of LLM Inference

### v4

@@ Abstract
This paper introduces a novel infrastructure-aware benchmarking framework for quantifying the environmental footprint of LLM inference across 30 state-of-the-art models as deployed in commercial data centers. Our framework combines public API performance data with region-specific environmental multipliers and statistical inference of hardware configurations. We additionally utilize cross-efficiency Data Envelopment Analysis (DEA) to rank models by performance relative to environmental cost. Our results show that o3 and DeepSeek-R1 emerge as the most energy-intensive models, consuming over 33 Wh per long prompt, more than 70 times the consumption of GPT-4.1 nano, and that Claude-3.7 Sonnet ranks highest in eco-efficiency. While a single short GPT-4o query consumes 0.42 Wh, scaling this to 700 million queries/day results in substantial annual environmental impacts. These include electricity use comparable to 35,000 U.S. homes, freshwater evaporation matching the annual drinking needs of 1.2 million people, and carbon emissions requiring a Chicago-sized forest to offset. These findings illustrate a growing paradox: Although AI is becoming cheaper and faster, its global adoption drives disproportionate resource consumption. Our study provides a standardized, empirically grounded methodology for benchmarking the sustainability of LLM deployments, laying a foundation for future environmental accountability in AI development and sustainability standards.
@@ 1 Introduction
However, the advancement of LLMs does involve shortcomings in environmental aspects. Training GPT-3 is estimated to consume 1,287 megawatt-hours (MWh) of electricity and emit over 550 metric tons of CO2e [12], while requiring more than 700 kiloliters (kL) of water for cooling alone [13], enough to fill two-thirds of an Olympic-sized swimming pool. Yet while training has been the focus of sustainability discussions, inference is emerging as the primary contributor to environmental costs. In contrast to training, which is conducted once or at intervals, inference occurs consistently and on a large scale. Recent estimates suggest inference can account for up to 90% of a model’s total lifecycle energy use [14, 15].
@@
Despite the growing environmental footprint of large-scale model deployment, a standard method to quantify the cost of inference at the prompt level remains absent. Existing frameworks [15, 19, 20] either lack the ability to benchmark proprietary models, the real-time granularity needed for deployment-specific prompt-level benchmarking, or are constrained to local setups, failing to capture the infrastructure complexity of production-scale inference. A core obstacle to developing more accurate assessments is the lack of information from commercial AI providers, as they do not disclose model-specific inference data, and existing environmental reports tend to aggregate emissions across entire cloud infrastructures without disaggregating by model or workload [17, 18]. This opacity hinders independent verification and undermines both scientific benchmarking and policy efforts aimed at regulating AI’s true environmental cost.
@@
To address these issues, we introduce a novel benchmarking framework to quantify the operational environmental footprint of LLM inference at the per-prompt level. Unlike existing studies [13, 15, 19], our method adopts a more comprehensive strategy by integrating performance metrics such as latency and throughput from public APIs with published GPU and system power specifications. Furthermore, we scale these combined data points using region-specific multipliers, including Power Usage Effectiveness (PUE) [21, 22], Water Usage Effectiveness (WUE) [21, 22], and Carbon Intensity Factors (CIF) [23, 24] to account for infrastructural overhead. This method enables us to evaluate the energy, water, and carbon effects of both open-source and proprietary models, a gap that, to our knowledge, has not been comprehensively explored in prior research. Additionally, we employ statistical analysis, including ANOVA and Tukey HSD, to estimate underlying hardware configurations. This framework’s effectiveness is shown by its use in more than 30 commercially deployed models and assessments across various infrastructure scenarios. Moreover, to contextualize resource use relative to model capability, we apply cross-efficiency Data Envelopment Analysis (DEA) to assess how effectively each model converts environmental inputs into performance. As a key application of this framework, we perform a case study to estimate the annual footprint of GPT-4o text generation based on scaled usage data. Our framework enables infrastructure-aware decision-making, empowers accountability, and provides a foundational step toward sustainability standards in AI deployment.
@@
The remainder of the paper is organized as follows. Section 2 reviews existing studies on the environmental impact of LLMs. Section 3 introduces key concepts, including hardware configurations and environmental multipliers. Section 4 details our framework for estimating inference-phase cost. Section 5 presents findings across 30 models. Section 6 provides a focused analysis of GPT-4o’s annual environmental footprint. Section 7 outlines key insights and implications. Section 8 summarizes the main takeaways and limitations and directions for future work.
@@ 2 Related Work
Most efforts focus on training and local model evaluation, lacking standardized, scalable methods, ignoring infrastructural overhead, and omitting resource categories such as water consumption and carbon emissions. Our work addresses these gaps by integrating API-based performance metrics with GPU and system power specifications and environmental multipliers to estimate the environmental impact of LLM inference at the prompt level in data centers. We infer deployment infrastructure through statistical analysis and apply DEA to contextualize environmental impact versus performance. Additionally, we conduct a case study estimating GPT-4o’s annual environmental footprint based on scaled usage data, providing the first infrastructure-aware, prompt-level benchmark of inference sustainability at scale.
@@ 4 Methodology
This section presents our novel methodology for estimating the environmental footprint of LLM inference. Our framework integrates model-specific performance metrics with infrastructure-level environmental multipliers to calculate operational energy consumption, water usage, and carbon emissions per query. We also evaluate eco-efficiency using DEA, mapping sustainability trade-offs against a composite performance benchmark.
@@ 4.1 Model Selection and Hardware Estimation
GPT-4.1 Apr, 2025 OpenAI Microsoft Azure DGX H200/H100 [35, 38] 10.20 [39] 1.12 [40] 0.30 [41] 3.142 [42] 0.3528 [37]
@@
GPT-4o mini July, 2024 OpenAI Microsoft Azure DGX A100* 6.50[43] 1.12 0.30 3.142 0.3528
@@
Claude-3.7 Sonnet Feb, 2025 Anthropic AWS DGX H200/H100 [47, 48] 10.20 1.14 [49] 0.18 [49] 3.142 0.385 [26]
Claude-3.7 Sonnet ET Feb, 2025
@@
Extended Thinking (ET).
We analyze 30 large language models across OpenAI, Anthropic, Meta, and DeepSeek. Table 1 summarizes each model’s deployment context, including provider, cloud host, hardware type and specifications, and regional environmental multipliers (PUE, WUE, CIF). All models are usually run on NVIDIA DGX systems using A100, H100, H200, or H800 GPUs [35, 31, 32, 33, 34]. U.S.-based providers such as OpenAI and Anthropic have acquired large volumes of H200 and H100 chips [38, 47, 48], making them the most probable choice for recent deployments. DeepSeek, which operates under U.S. export restrictions, uses the H800, NVIDIA’s export-compliant GPU for the Chinese market [44, 50]. Both the H200 and H800 retain the same Hopper architecture and peak power draw as the H100, with system-level energy characteristics that are nearly identical [51]. While the H200 achieves greater energy efficiency due to faster memory and higher bandwidth, and the H800 may exhibit reduced performance due to export-related firmware limitations, both maintain the same peak power draw, thermal design profile, and system-level utilization characteristics as the H100 [44, 51]. These architectural differences affect throughput and latency, resulting in higher or lower energy consumed per token, but do not impact total system power demand under load. We therefore treat H100, H200, and H800 as equivalent in our power modeling, since our estimates are based on power draw and utilization rather than task-level performance.
Environmental multipliers such as PUE, WUE, and CIF are assigned based on the cloud provider and regional deployment environments. OpenAI models, hosted on Microsoft Azure, use Azure-reported PUE, site-level WUE, and CIF values, supplemented by U.S. national averages for source WUE. For AWS-hosted Anthropic and Meta models, we apply AWS-reported PUE and site-level WUE, using U.S. national averages for source WUE and CIF due to limited public disclosures. For DeepSeek, we use Chinese data centers’ national averages for all multipliers.
@@ 4.2 Per-Query Energy Consumption Estimation
To estimate the per-query energy consumption, we introduce a formula that serves as the core of our infrastructure-aware framework. This model integrates performance data [52], which evaluates LLMs under standardized conditions. For each model, we extract latency and tokens-per-second (TPS) across three prompt configurations that represent real-world use: Short-form (100 input, 300 output tokens), medium (1,000 input, 1,000 output), and long-form (10,000 input, 1,500 output). Metrics are reported as distributions over the 5th, 25th, 50th (median), 75th, and 95th percentiles, reflecting variability across multiple test runs for each model and prompt configuration. We compute runtime estimates by pairing available latency and TPS quantiles to construct a joint distribution that captures decoding throughput and fixed overhead variability. This modeling ensures that downstream energy calculations reflect the stochastic nature of inference workloads. Our model computes the per-query energy as:
@@
, where $P_{\text{GPU}}$ and $P_{\text{non-GPU}}$ represent the maximum rated GPU and non-GPU system power, respectively, measured at the node level. $U_{\text{GPU}}$ denotes the aggregate GPU power draw, incorporating both the number of GPUs assigned and their per-GPU load, while $U_{\text{non-GPU}}$ similarly reflects the aggregate power utilization of non-GPU components. Latency refers to the time to first token generation, which corresponds to the time required to process the input prompt. TPS represents the generation rate of output tokens; therefore, dividing the output length by TPS yields the time required to generate the response. The formula is evaluated across all quantile pairs and output lengths to produce a range of energy estimates per model configuration and output size, incorporating the PUE factor to account for data center-level overheads.
@@ 4.3.1 GPT-4, GPT-4 Turbo, and GPT-4o mini Hardware Estimation
Provider (System) Provider (System) 300 tokens 1000 tokens 1500 tokens
@@ 5 Experimental Evaluation
We benchmark the environmental footprint of 30 LLMs across three modalities: Energy consumption, water usage, and carbon emissions, based on equations 1, 3, and 4, respectively. For the long-form query evaluation, GPT-4 and LLaMA-3 (8B and 70B) are excluded due to context window limitations.
@@ 5.1 Energy Consumption
GPT-4.1 0.918 ± 0.498 2.513 ± 1.286 4.233 ± 1.968
GPT-4.1 mini 0.421 ± 0.197 0.847 ± 0.379 1.590 ± 0.801
GPT-4.1 nano 0.103 ± 0.037 0.271 ± 0.087 0.454 ± 0.208
o4-mini (high) 2.916 ± 1.605 5.039 ± 2.764 5.666 ± 2.118
GPT-4.5 6.723 ± 1.207 20.500 ± 3.821 30.495 ± 5.424
o3 7.026 ± 3.663 21.414 ± 14.273 39.223 ± 20.317
o3-mini (high) 2.319 ± 0.670 5.128 ± 1.599 4.596 ± 1.453
o3-mini 0.850 ± 0.336 2.447 ± 0.943 2.920 ± 0.684
o1 4.446 ± 1.779 12.100 ± 3.922 17.486 ± 7.701
o1-mini 0.631 ± 0.205 1.598 ± 0.528 3.605 ± 0.904
GPT-4o (Mar ’25) 0.421 ± 0.127 1.214 ± 0.391 1.788 ± 0.363
GPT-4o mini 0.421 ± 0.082 1.418 ± 0.332 2.106 ± 0.477
GPT-4 Turbo 1.656 ± 0.389 6.758 ± 2.928 9.726 ± 2.686
GPT-4 1.978 ± 0.419 6.512 ± 1.501
DeepSeek-R1 23.815 ± 2.160 29.000 ± 3.069 33.634 ± 3.798
DeepSeek-V3 3.514 ± 0.482 9.129 ± 1.294 13.838 ± 1.797
Claude-3.7 Sonnet 0.836 ± 0.102 2.781 ± 0.277 5.518 ± 0.751
Claude-3.7 Sonnet ET 3.490 ± 0.304 5.683 ± 0.508 17.045 ± 4.400
LLaMA-3-8B 0.092 ± 0.014 0.289 ± 0.045
LLaMA-3-70B 0.636 ± 0.080 2.105 ± 0.255
LLaMA-3.1-8B 0.103 ± 0.016 0.329 ± 0.051 0.603 ± 0.094
LLaMA-3.1-70B 1.101 ± 0.132 3.558 ± 0.423 11.628 ± 1.385
LLaMA-3.1-405B 1.991 ± 0.315 6.911 ± 0.769 20.757 ± 1.796
LLaMA-3.2 1B 0.070 ± 0.011 0.218 ± 0.035 0.342 ± 0.056
LLaMA-3.2 3B 0.115 ± 0.019 0.377 ± 0.066 0.573 ± 0.098
LLaMA-3.2-vision 11B 0.071 ± 0.011 0.214 ± 0.033 0.938 ± 0.163
LLaMA-3.2-vision 90B 1.077 ± 0.096 3.447 ± 0.302 5.470 ± 0.493
LLaMA-3.3 70B 0.247 ± 0.032 0.857 ± 0.113 1.646 ± 0.220
Figure 2 and Table 4 highlight how energy consumption scales with prompt length and model architecture, revealing substantial disparities across systems. GPT-4.1 nano remains the most efficient overall, requiring only 0.454 Wh for long prompts (approximately 7,000 words of input and 1,000 words of output). In contrast, o3 consumes 39.223 Wh, while DeepSeek-R1 and GPT-4.5 consume 33.634 Wh and 30.495 Wh, respectively, which is over seventy times the energy use of GPT-4.1 nano. To contextualize, a single long query to o3 or DeepSeek-R1 may consume as much electricity as running a 65-inch LED television ($\approx 130\,\mathrm{W}$) for roughly 20–30 minutes. Although o3 and DeepSeek-R1 rely heavily on chain-of-thought prompting, GPT-4.5 stands out for its relatively high energy use, despite not being a multi-step reasoning model. This suggests inefficiencies rooted in model architecture.
Claude-3.7 Sonnet ET presents a notable exception. While it supports chain-of-thought reasoning, it consumes only 17.045 Wh for long-form input, which is less than half the energy of o3. Similarly, GPT-4o, OpenAI’s current default model, demonstrates strong energy efficiency, requiring just 1.788 Wh for long prompts and 0.42 Wh for short ones. Interestingly, GPT-4o mini, although substantially smaller in parameter count, consumes slightly more energy per query than GPT-4o due to its deployment on less efficient A100 hardware instead of H100s or H200s, illustrating that deployment infrastructure can overshadow model size in determining real-world energy use.

### v5

@@ Abstract
[…]
1 University of Rhode Island  2 University of Tunis  3 Providence College
Live Dashboard: Power BI Dashboard
@@ 1 Introduction
However, the advancement of LLMs does involve shortcomings in environmental aspects. Training GPT-3 is estimated to consume 1,287 megawatt-hours (MWh) of electricity and emit over 550 metric tons of CO2 equivalent ($\text{CO}_{2}\text{e}$) [12], while requiring more than 700 kiloliters (kL) of water for cooling alone [13], enough to fill two-thirds of an Olympic-sized swimming pool. Yet while training has been the focus of sustainability discussions, inference is emerging as the primary contributor to environmental costs. In contrast to training, which is conducted once or at intervals, inference occurs consistently and on a large scale. Recent estimates suggest inference can account for up to 90% of a model’s total lifecycle energy use [14, 15].
@@
Despite the growing environmental footprint of large-scale model deployment, a standard method to quantify the cost of inference at the prompt level remains absent. A core obstacle to developing more accurate assessments is the lack of model-specific inference data for commercial AI models. Existing environmental reports tend to aggregate emissions across entire cloud infrastructures without disaggregating by model or workload [16, 17]. This lack of public information hinders independent verification and undermines both scientific benchmarking and policy efforts aimed at regulating AI’s true environmental cost.
@@
To address these issues, we introduce a novel infrastructure-aware benchmarking framework to quantify the operational environmental footprint of LLM inference at the per-prompt level as deployed in data centers. Unlike existing studies [13, 15, 18], our method adopts a more comprehensive strategy by integrating performance metrics such as latency and throughput from public APIs with published GPU and system power specifications. Furthermore, we scale these combined data points using company-specific multipliers, including Power Usage Effectiveness (PUE) [19, 20], Water Usage Effectiveness (WUE) [19, 20], and Carbon Intensity Factors (CIF) [21, 22] to account for infrastructural overhead. This method enables us to evaluate the energy, water, and carbon effects of both open-source and proprietary models, a gap that, to our knowledge, has not been comprehensively explored in prior research. Additionally, we employ statistical analysis, including ANOVA and Tukey HSD, to estimate underlying hardware configurations. To enhance transparency and reproducibility, we also developed an automated and interactive Power BI dashboard that visualizes the daily fluctuations in the energy, water, and carbon footprint of an extended list of models across multiple data centers. This novel dashboard incorporates new models as they get released. Moreover, to contextualize resource use relative to model capability, we apply cross-efficiency Data Envelopment Analysis (DEA) to assess how effectively each model converts environmental inputs into performance. As a key application of this framework, we perform a case study to estimate the footprint of GPT-4o text generation based on scaled usage data. We further extend our analysis to GPT-5, focusing on the disparities in energy consumption between queries that involve different levels of reasoning. Our framework enables infrastructure-aware decision-making, empowers accountability, and provides a foundational step toward sustainability standards in AI deployment.
@@
The remainder of the paper is organized as follows. Section 2 reviews existing studies on the environmental impact of LLMs. Section 3 introduces key concepts, including hardware configurations and environmental multipliers. Section 4 details our framework for estimating inference-phase cost. Section 5 presents findings across 30 models. Section 6 provides a focused analysis of GPT-4o’s annual environmental footprint and section 7 analyzes the impact of GPT-5’s adapative model routing. Section 8 outlines key insights and implications. Section 9 summarizes the main takeaways and limitations and directions for future work.
@@ 2 Related Work
Most efforts focus on training and local model evaluation, lacking standardized, scalable methods, ignoring infrastructural overhead, and omitting resource categories such as water consumption and carbon emissions. Our work addresses these gaps by integrating API-based performance metrics with GPU and system power specifications and environmental multipliers to estimate the environmental impact of LLM inference at the prompt level in data centers. We infer deployment infrastructure through statistical analysis and apply DEA to contextualize environmental impact versus performance. Additionally, we conduct two case studies estimating GPT-4o’s annual environmental footprint based on scaled usage data and analyzing the impact of GPT-5’s adapative model routing, providing the first infrastructure-aware, prompt-level benchmark of inference sustainability at scale.
@@ 4 Methodology
This section presents our novel methodology for estimating the environmental footprint of LLM inference. Our framework integrates model-specific performance metrics with infrastructure-level environmental multipliers to calculate operational energy consumption, water usage, and carbon emissions per query. We also evaluate eco-efficiency using DEA, mapping sustainability trade-offs against a composite performance benchmark, and develop an interactive dashboard for a more thorough analysis.
@@ 4.1 Model Selection and Hardware Estimation
GPT-4.1 Apr, 2025 OpenAI Microsoft Azure DGX H200/H100 [30, 31] 10.20 [32] 1.12 [33] 0.30 [34] 4.35 [35] 0.35 [36]
@@
GPT-4o mini July, 2024 OpenAI Microsoft Azure DGX A100* 6.50[37] 1.12 0.30 4.35 0.35
@@
DeepSeek-R1 Jan, 2025 Deepseek Microsoft Azure DGX H200/H100 10.20 1.12 0.30 4.35 0.35
DeepSeek-V3 Dec, 2024
Claude-3.7 Sonnet Feb, 2025 Anthropic AWS DGX H200/H100 [41, 42] 10.20 1.14 [43] 0.18 [43] 5.11 [35] 0.287 [44]
@@
We analyze 30 large language models across OpenAI, Anthropic, Meta, and DeepSeek. Table 1 summarizes each model’s deployment context, including provider, cloud host, hardware type and specifications, and company-specific environmental multipliers (PUE, WUE, CIF). All models are usually run on NVIDIA DGX systems using A100, H100, H200, or H800 GPUs [30, 45, 46, 47, 48]. U.S.-based providers such as OpenAI and Anthropic have acquired large volumes of H200 and H100 chips [31, 41, 42], making them the most probable choice for recent deployments. DeepSeek, which operates under U.S. export restrictions, uses the H800, NVIDIA’s export-compliant GPU for the Chinese market [38, 49]. Both the H200 and H800 retain the same Hopper architecture and peak power draw as the H100, with system-level energy characteristics that are nearly identical [50]. While the H200 achieves greater energy efficiency due to faster memory and higher bandwidth, and the H800 may exhibit reduced performance due to export-related firmware limitations, both maintain the same peak power draw, thermal design profile, and system-level utilization characteristics as the H100 [38, 50]. These architectural differences affect throughput and latency, resulting in higher or lower energy consumed per token, but do not impact total system power demand under load. We therefore treat H100, H200, and H800 as equivalent in our power modeling, since our estimates are based on power draw and utilization rather than task-level performance.
Environmental multipliers such as PUE, WUE, and CIF are assigned according to each cloud provider’s data center locations and corresponding regional grid characteristics. For OpenAI and DeepSeek models hosted on Microsoft Azure, we use Azure-reported PUE and site-level WUE values, while CIF and source-level WUE are derived from the specific geographic locations of Microsoft data centers around the world. For AWS-hosted models, including those from Anthropic and Meta, we apply AWS-reported PUE and site-level WUE, and compute CIF and source-level WUE based on the regional distribution of AWS data centers used for inference. For DeepSeek models that are deployed in Chinese datacenters, we adopt the average PUE and site-level WUE of the thirty most efficient data centers in China, while CIF and source-level WUE are determined using the regional locations of its known or reported data center deployments.
@@ 4.2 Per-Query Energy Consumption Estimation
To quantify the energy required for a single inference, we introduce a probabilistic framework that captures the stochastic nature of LLM workloads. The model integrates standardized performance data [51], which report latency to first-token generation ($L$) and tokens-per-second (TPS, denoted $R$) across empirical quantiles (5th, 25th, 50th, 75th, and 95th percentiles) and three representative prompt configurations: short-form (100 input, 300 output tokens), medium (1,000 input, 1,000 output), and long-form (10,000 input, 1,500 output), reflecting variability across multiple test runs for each model and prompt configuration.
To model realistic runtime behavior, we construct a joint distribution of $L$ and $R$ using a Gaussian copula with correlation coefficient $\rho=-0.3$, capturing the negative dependence typically observed between latency and TPS. From this distribution, we draw 10,000 correlated samples $(L_{i},R_{i})$, each representing one plausible inference scenario. The culmination of this infrastructure-aware framework is the introduction of our novel formula to precisely estimate the per-query energy consumption:
Let $L_{i}$ captures the initialization latency and $\tfrac{\text{Output Length}}{R_{i}}$ represents the time it takes to generate the response. Also, let $P_{\text{GPU}}$ and $P_{\text{non-GPU}}$ denote the rated power draw (in kW) of the GPU subsystem and the non-GPU subsystem (e.g., CPUs, SSDs, network, and cooling control electronics), respectively. The parameters $U_{\text{GPU,min}}$ and $U_{\text{GPU,max}}$ represent the minimum and maximum GPU utilization fractions observed during inference, while $U_{\text{non-GPU}}$ represents the average utilization fraction for non-GPU components. PUE factor is also incorporated to account for datacenter-level overheads.
We compute energy consumption at the lower and upper utilization bounds as:
@@
We also define an expected per-query energy as a weighted combination of both scenarios ($w_{\max}=0.5$), and the framework aggregates all Monte Carlo draws to produce a distribution of per-query energy outcomes. The final metrics are reported as the sample mean and standard deviation:
$E_{i,\text{exp}}=w_{\max}E_{i,\max}+(1-w_{\max})E_{i,\min},\quad\bar{E}_{\text{query}}=\mathbb{E}[E_{i,\text{exp}}],\quad\sigma_{E_{\text{query}}}=\sqrt{\text{Var}[E_{i,\text{exp}}]}$ (2)
This stochastic formulation captures variability in runtime, hardware utilization, and data-center efficiency, enabling robust and reproducible estimation of per-query energy consumption across diverse inference conditions.
@@ 4.3.1 GPT-4, GPT-4 Turbo, and GPT-4o mini Hardware Estimation
Group 1 Group 2 300 tokens 1000 tokens 1500 tokens
@@ 4.5 Eco-Efficiency via Data Envelopment Analysis (DEA)
4.6 Power BI Dashboard
To democratize access to these novel assessments, we built and deployed an automated Power BI dashboard that runs our entire framework in real time, a first-of-its-kind tool for continuously tracking AI inference sustainability. The data are scraped daily from the Artificial Analysis website, cleaned automatically, and then visualized on Power BI as seen in Figures 2(a) and 2(b). The main dashboard displays the average and standard deviation of energy use, water consumption (site, source, and combined), and carbon emissions for the three query sizes. It also visualizes latency and TPS fluctuations, benchmark results, and the total environmental impact when scaling up to $1$, 5050, or $100$ billion queries, compared with real-world equivalents such as household electricity use, annual drinking needs, and transportation emissions. Users can filter by company, model size, query size, or sustainability metric, and download the full dataset. Additionally, the dashboard tracks day-to-day changes in each model’s footprint, visualizing time-series trends and the average in energy, water, and carbon metrics across data centers and hardware setups. It includes an extended list of models beyond those analyzed in this study and automatically incorporates new ones as they are released, allowing continuous monitoring of inference-phase sustainability and cross-model comparisons over time.
(a) Overview of the main dashboard displaying the energy consumption per model, latency, TPS, benchmark scores, and equivalent environmental impacts for an example model (GPT-5 minimal).
(b) Overview of the timeseries dashboard displaying average energy consumption per model, and the daily fluctuations of the selected model (Grok 4).
Figure 2: Visual overview of the AI sustainability dashboard.
@@ 5 Experimental Evaluation
We benchmark the environmental footprint of 30 LLMs across three modalities: Energy consumption, water usage, and carbon emissions, based on equations 2, 4, and 5, respectively. For the long-form query evaluation, GPT-4 and LLaMA-3 (8B and 70B) are excluded due to context window limitations.
@@ 5.1 Energy Consumption
GPT-4.1 0.871 ± 0.302 3.161 ± 0515 4.833 ± 0.650
GPT-4.1 mini 0.450 ± 0.081 1.545 ± 0.211 2.122 ± 0.348
GPT-4.1 nano 0.207 ± 0.047 0.575 ± 0.108 0.827 ± 0.094
o4-mini (high) 3.649 ± 1.468 7.380 ± 2.177 7.237 ± 1.674
o3 1.177 ± 0.224 5.153 ± 2.107 12.222 ± 1.082
o3-mini (high) 3.012 ± 0.991 6.865 ± 1.33 5.389 ± 1.183
o3-mini 0.674 ± 0.015 2.423 ± 0.237 3.525 ± 0.168
o1 2.268 ± 0.654 4.047 ± 0.497 6.181 ± 0.877
o1-mini 0.535 ± 0.182 1.547 ± 0.405 2.317 ± 0.530
GPT-4o (Mar ’25) 0.423 ± 0.085 1.215 ± 0.241 2.875 ± 0.421
GPT-4o mini 0.577 ± 0.139 1.897 ± 0.570 3.098 ± 0.639
GPT-4 Turbo 1.699 ± 0.355 5.940 ± 1.441 9.877 ± 1.304
GPT-4 1.797 ± 0.259 6.925 ± 1.553
DeepSeek-R1 (DS) * 19.251 ± 9.449 24.596 ± 9.4 29.078 ± 9.725
DeepSeek-V3 (DS) * 2.777 ± 0.223 8.864 ± 0.724 13.162 ± 1.126
DeepSeek-R1 (AZ) 2.353 ± 1.129 4.331 ± 1.695 7.410 ± 2.159
DeepSeek-V3 (AZ) 0.742 ± 0.125 2.165 ± 0.578 3.696 ± 0.221
Claude-3.7 Sonnet 0.950 ± 0.040 2.989 ± 0.201 5.671 ± 0.302
Claude-3.5 Sonnet 0.973 ± 0.066 3.638 ± 0.256 7.772 ± 0.345
Claude-3.5 Haiku 0.975 ± 0.063 4.464 ± 0.283 8.010 ± 0.338
LLaMA-3-8B 0.108 ± 0.002 0.370 ± 0.005
LLaMA-3-70B 0.861 ± 0.022 2.871 ± 0.094
LLaMA-3.1-8B 0.052 ± 0.008 0.172 ± 0.015 0.443 ± 0.028
LLaMA-3.1-70B 1.271 ± 0.020 4.525 ± 0.053 19.183 ± 0.560
LLaMA-3.1-405B 2.226 ± 0.142 9.042 ± 0.385 25.202 ± 0.526
LLaMA-3.2 1B 0.109 ± 0.013 0.342 ± 0.025 0.552 ± 0.059
LLaMA-3.2 3B 0.143 ± 0.006 0.479 ± 0.017 0.707 ± 0.020
LLaMA-3.2-vision 11B 0.078 ± 0.021 0.242 ± 0.071 1.087 ± 0.060
LLaMA-3.2-vision 90B 1.235 ± 0.054 4.534 ± 0.448 6.852 ± 0.780
LLaMA-3.3 70B 0.237 ± 0.023 0.760 ± 0.079 1.447 ± 0.188
* DeepSeek Host
Microsoft Azure Host
Figure 3 and Table 4 highlight how energy consumption scales with prompt length and model architecture, revealing wide disparities across systems. LLaMA-3.1-8B is the most efficient, requiring only 0.443 Wh for long prompts (approximately 7,000 words of input and 1,000 words of output), followed by LLaMA-3.2 1B and LLaMA-3.2 3B at 0.552 Wh and 0.707 Wh, respectively. GPT-4.1 nano remains among the most efficient proprietary models at 0.827 Wh, but still consumes nearly twice the energy of LLaMA-3.1-8B. In contrast, DeepSeek-R1 (DS) consumes 29.075 Wh, around sixty five times more than the most efficient model, underscoring the large overhead of reasoning models.
The LLaMA family shows clear scaling effects: energy use rises from 0.443 Wh at 8B parameters to 25.202 Wh at 405B, illustrating steep power demands at high parameter counts. Additionally, the DeepSeek models reveal striking infrastructure effects. DeepSeek-R1 and DeepSeek-V3 hosted on DeepSeek’s own servers consume 29.078 Wh and 13.162 Wh, while the same models on Azure use just 7.410 Wh and 3.696 Wh, over 70% less energy. This gap highlights that hardware and data center efficiency, not model design alone, drives real-world energy use. For context, a single long query to DeepSeek-R1 (DS) consumes about as much electricity as running a 65-inch LED television ($\approx 130\,\mathrm{W}$) for roughly 13 minutes. GPT-4o and GPT-4o mini also show that infrastructure can outweigh model size in determining energy efficiency. For instance GPT-4o consumes around 2.875 Wh while GPT-4o mini’s consumption is slightly higher at 3.098 Wh due to deployment on A100 hardware instead of H100s.